2
5a + 13a + 8

factor completely
1
Expert's answer
2012-02-14T10:29:22-0500
Let's factor the expression by separating the complete square:
5a^2+13a+8=5(a^2+2a+1)+3a+3=5(a+1)^2+3(a+1)=(a+1)(5(a+1)+3)=(a+1)(5a+5+3)=(a+1)(5a+8)
Also we can do it in another way by finding roots of the given polynome:
5a^2+13a+8=0
D=13^2-4*8*5=169-160=9
a1=(-13+3)/10=-1
a2=(-13-3)/10=-16/10
Now we can factor the polynome as following:
5a^2+13a+8=5(a+1)(a+16/10)=(a+1)(5a+8) - no surprises, we've obtained the same result

For our visitors we also created video where factorisation of quadratic polynom is considered. Please take a look!